# LlamaFactory: Unified Efficient Fine-Tuning of 100+ Language Models

### v2

@@ Abstract
Efficient fine-tuning is vital for adapting large language models (LLMs) to downstream tasks. However, it requires non-trivial efforts to implement these methods on different models. We present LlamaFactory, a unified framework that integrates a suite of cutting-edge efficient training methods. It allows users to flexibly customize the fine-tuning of 100+ LLMs without the need for coding through the built-in web UI LlamaBoard. We empirically validate the efficiency and effectiveness of our framework on language modeling and text generation tasks. It has been released at https://github.com/hiyouga/LLaMA-Factory and already received over 13,000 stars and 1,600 forks.
@@
Yaowei Zheng1, Richong Zhang1, Junhao Zhang1, Yanhan Ye1, Zheyan Luo1, Yongqiang Ma2 1School of Computer Science and Engineering, Beihang University, China 2School of Software and Microelectronics, Peking University, China hiyouga@buaa.edu.cn, zhangrc@act.buaa.edu.cn, {zhang.jh,yeyanhan,akamya}@buaa.edu.cn, codingma@pku.edu.cn Demonstration video: https://youtu.be/W29FgeZEpus
@@ 1 Introduction
Fine-tuning extremely large number of parameters with limited resources becomes the main challenge of adapting LLM to downstream tasks. A popular solution is efficient fine-tuning (Houlsby et al., 2019; Hu et al., 2022; Ben Zaken et al., 2022; Dettmers et al., 2023; Zhao et al., 2024), which reduces the training cost of LLMs when adapting to various tasks. However, the community contributes various methods for efficient fine-tuning LLMs, lacking a systematic framework that adapts and unifies these methods to different LLMs and provides a friendly interface for user customization.
@@
LlamaFactory is open-sourced under the Apache-2.0 license. It has already garnered over 13,000 stars and 1,600 forks on the GitHub [^1] https://github.com/hiyouga/LLaMA-Factory, and hundreds of open-source models have been built upon LlamaFactory on the Hugging Face Hub [^2] https://huggingface.co/models?other=llama-factory. For example, the well-known GemSUra-7B (Nguyen et al., 2024) is built based on LlamaFactory, which first reveals the cross-lingual abilities of Gemma (Mesnard et al., 2024). Furthermore, dozens of research studies utilize our framework to explore new methods in LLMs, such as Wang et al. (2023a); Yu et al. (2023); Bhardwaj et al. (2024).
@@ 2 Efficient Fine-Tuning Techniques
Efficient LLM fine-tuning techniques can be divided into two main categories: those focused on optimization and those aimed at computation. The primary objective of efficient optimization techniques is to adjust the parameters of LLMs while keeping costs to a minimum. On the other hand, efficient computation methods seek to decrease the time or space for the required computation in LLMs. The methods included in LlamaFactory are listed in Table 1. We will present these efficient fine-tuning techniques and show the substantial efficiency improvement achieved by incorporating them into our framework in the following sections.
Table 1: Efficient fine-tuning techniques featured in LlamaFactory. Techniques that are compatible with each other are marked with ✓, while those that are not compatible are marked with ✗.
Freeze-tuning GaLore LoRA DoRA
@@ 2.1 Efficient Optimization
Firstly, we provide an overview of the efficient optimization techniques utilized in LlamaFactory. The freeze-tuning method (Houlsby et al., 2019) involves freezing a majority of parameters while fine-tuning the remaining parameters in a small subset of decoder layers. Another method called gradient low-rank projection (GaLore) (Zhao et al., 2024) projects gradients into a lower-dimensional space, facilitating full-parameter learning in a memory-efficient manner. On the contrary, the low-rank adaptation (LoRA) (Hu et al., 2022) method freezes all pre-trained weights and introduces a pair of trainable low-rank matrices to the designated layer. When combined with quantization, this approach is referred to as QLoRA (Dettmers et al., 2023), which additionally reduces the memory usage. The weight-decomposed low-rank adaptation (DoRA) (Liu et al., 2024) method breaks down pre-trained weights into magnitude and direction components, applying LoRA solely on the directional components to enhance the fine-tuning of LLMs. LoRA+ (Hayou et al., 2024) is proposed to overcome the sub-optimality of LoRA.
@@ 2.2 Efficient Computation
In LlamaFactory, we integrate a range of efficient computational techniques. Commonly utilized techniques encompass mixed precision training (Micikevicius et al., 2018) and activation checkpointing (Chen et al., 2016). Drawing insights from the examination of the input-output (IO) expenses of the attention layer, flash attention (Dao et al., 2022) introduces a hardware-friendly approach to enhance attention computation. S2 attention (Chen et al., 2024b) tackles the challenge of extending context in block-sparse attention, thereby diminishing memory usage in fine-tuning long-context LLMs. Various quantization strategies (Dettmers et al., 2022a; Frantar et al., 2023; Lin et al., 2023; Egiazarian et al., 2024) decrease memory requirements in large language models (LLMs) by utilizing lower-precision representations for weights. Nevertheless, the fine-tuning of quantized models is restricted to the adapter-based techniques like LoRA (Hu et al., 2022). Unsloth (Han and Han, 2023) incorporates Triton (Tillet et al., 2019) for implementing the backward propagation of LoRA, which reduces floating-point operations (FLOPs) during gradient descent and leads to expedited LoRA training.
LlamaFactory effectively combines these techniques into a cohesive structure that greatly enhances the efficiency of LLM fine-tuning. This results in a reduction of the memory footprint from 18 bytes per parameter during mixed precision training (Micikevicius et al., 2018) or 8 bytes per parameter in bfloat16 training (Le Scao et al., 2022) to only 0.6 bytes per parameter. Further elaboration on our framework will be provided in the subsequent section.
@@ 3 LlamaFactory Framework
LlamaFactory consists of three main modules: Model Loader, Data Worker, and Trainer. The Model Loader prepares various architectures for fine-tuning, which supports over 100 LLMs. The Data Worker processes data from different tasks through a well-designed pipeline, which support more than 50 datasets. The Trainer unifies efficient fine-tuning methods to adapt these models on different tasks and datasets, which offers four training approaches. LlamaBoard provides a friendly visual interface for the above modules, enabling users to configure and launch individual LLM fine-tuning process in a codeless way and monitor the training status on the fly. We illustrate the overall architecture of LlamaFactory in Figure 1.
LlamaBoardArgument ConfiguratorTraining Status MonitorTrainerOptimizationApproachesLoRA+GaLoreOffloadPartitionPre-trainSFTRLHFDPOModel LoaderData WorkerInitializationPatchesQuantizationAdaptersAligningLoadingPreprocessMerging100+ Pre-Trained Models50+ NLP Datasets
@@ 3.1 Model Loader
This section initially presents the four components in Model Loader: Model Initialization, Model Patching, Model Quantization, and Adapter Attaching, followed by a description of our approach of adapting to a wide range of devices by handling the parameter float-point precision when fine-tuning.
@@
We employ the AutoModel API of Transformers (Wolf et al., 2020) for loading models and initializing parameters. In order to make our framework compatible with different model architectures, we establish a model registry to store the type of each layer, thereby facilitating the utilization of efficient fine-tuning techniques more straightforward. In cases where the vocabulary size of the tokenizer exceeds the capacity of the embedding layer, we resize the layer and initialize new parameters with noisy mean initialization. To determine the scaling factor for RoPE scaling (Chen et al., 2023), we compute it as the ratio of the maximum input sequence length to the context length of the model.
@@
To enable flash attention and S2 attention, we employ a monkey patch to replace the forward computation of models. Nevertheless, since flash attention has been supported since Transformers 4.34.0, we use the API to enable flash attention. To prevent excessive partitioning of the dynamic modules, we set the mixture-of-experts (MoE) blocks as leaf modules when we optimize it under DeepSpeed ZeRO-3 (Rasley et al., 2020).
@@
We automatically identify the appropriate layers using the model registry to attach adapters. Adapters are attached to a subset of layers by default to save memory, but attaching them to all linear layers may yield better performance (Dettmers et al., 2023). The PEFT (Mangrulkar et al., 2022) library offers an extremely convenient way to attach adapters such as LoRA (Hu et al., 2022), rsLoRA (Kalajdzievski, 2023), and DoRA (Liu et al., 2024). We replace the backward computation with the one of Unsloth (Han and Han, 2023) to accelerate LoRA. To perform reinforcement learning from human feedback (RLHF), we add a value head upon the model, a linear layer that maps the representation of each token to a scalar.
@@
We handle the floating-point precision of pre-trained models based on the capabilities of devices. For NVIDIA GPUs, we use bfloat16 precision if the computation capability is 8.0 or higher. Otherwise, float16 is adopted. We use float16 for Ascend NPUs and AMD GPUs and float32 for non-CUDA devices. Note that loading bfloat16 models with float16 precision may lead to overflow issues. In mixed precision training, we set all trainable parameters to float32. Nevertheless, we retain the trainable parameters as bfloat16 in bfloat16 training.
@@ 3.2 Data Worker
"system": "…", "tools": "…"}
@@
LlamaFactory is designed for fine-tuning the text generative models, which is primarily used in chat completion. Chat template is a crucial component in these models, because it is highly related to the instruction-following abilities of these models. Therefore, we provide dozens of chat templates that can be automatically chosen according to the model type. We encode the sentence after applying the chat template using the tokenizer. By default, we only compute loss on the completions, while the prompts are disregarded (Taori et al., 2023). Optionally, we can utilize sequence packing (Krell et al., 2021) to reduce the training time, which is automatically enabled when performing generative pre-training. Appendix C shows details of our template design.
@@ 3.3 Trainer
We integrate state-of-the-art efficient fine-tuning methods, including LoRA+ (Hayou et al., 2024) and GaLore (Zhao et al., 2024) to the trainer by replacing the default component. These training approaches are independent of the trainer, making them easily applicable to various tasks. We utilize the trainer of Transformers (Wolf et al., 2020) for pre-training and SFT, while adopting the trainers of TRL (von Werra et al., 2020) for RLHF and DPO. The tailored data collators are leveraged to differentiate trainers of various training approaches. To match the input format of the trainer for preference data, we build $2n$ samples in a batch where the first $n$ samples are chosen examples and the last $n$ samples are rejected examples.
@@
Allowing RLHF training on consumer devices is a useful property for LLM fine-tuning. However, it is difficult because RLHF training requires four different models. To address this problem, we propose model-sharing RLHF, enabling entire RLHF training with no more than one pre-trained model. Concretely, we first train an adapter and a value head with the objective function for reward modeling, allowing the model to compute reward scores. Then we initialize another adapter and value head and train them with the PPO algorithm (Ouyang et al., 2022). The adapters and value heads are dynamically switched through the set_adapter and disable_adapter APIs of PEFT (Mangrulkar et al., 2022) during training, allowing the pre-trained model to serve as policy, value, reference, and reward models simultaneously. To the best of our knowledge, this is the first method that supports RLHF training on consumer devices.
@@
We can combine the above trainers with DeepSpeed (Rasley et al., 2020) for distributed training. Leveraging the DeepSpeed ZeRO optimizer, the memory consumption can be further reduced via partitioning or offloading.
@@ 3.4 Utilities
Comprehensive Evaluation
We include several metrics for evaluating LLMs, including multiple-choice tasks such as MMLU (Hendrycks et al., 2021), CMMLU (Li et al., 2023a), and C-Eval (Huang et al., 2023), as well as calculating text similarity scores like BLEU-4 (Papineni et al., 2002) and ROUGE (Lin, 2004).
@@ 3.5 LlamaBoard: A Unified Interface for LlamaFactory
LlamaBoard is a unified user interface based on Gradio (Abid et al., 2019) that allows users to customize the fine-tuning of LLMs without writing any code. It offers a streamlined model fine-tuning and inference service, enabling users to easily leveraging 100+ LLMs and 50+ datasets in their practice. LlamaBoard has the following notable features:
Easy Configuration LlamaBoard allows users to customize fine-tuning arguments through interaction with the web interface. We provide default values for many arguments that are recommended for most users, simplifying the configuration process. Moreover, users can preview datasets on the web UI to check their customized format.
Monitorable Training During the training process, the training logs and loss curves are visualized and updated in real time, allowing users to monitor the training progress. This feature provides valuable insights to analyze the fine-tuning process.
Flexible Evaluation LlamaBoard supports calculating the text similarity scores on the datasets to automatically evaluate the model or performing human evaluation by chatting with the model.
Multilingual Support LlamaBoard provides localization files, facilitating the integration of new languages for rendering the interface. Currently we support three languages: English, Russian and Chinese, which allows a broader range of users to utilize LlamaBoard for fine-tuning LLMs.

### v3

@@ Abstract
Efficient fine-tuning is vital for adapting large language models (LLMs) to downstream tasks. However, it requires non-trivial efforts to implement these methods on different models. We present LlamaFactory, a unified framework that integrates a suite of cutting-edge efficient training methods. It provides a solution for flexibly customizing the fine-tuning of 100+ LLMs without the need for coding through the built-in web UI LlamaBoard. We empirically validate the efficiency and effectiveness of our framework on language modeling and text generation tasks. It has been released at https://github.com/hiyouga/LLaMA-Factory and received over 24,000 stars and 3,000 forks.
@@
Yaowei Zheng1, Richong Zhang1thanks: Corresponding author, Junhao Zhang1, Yanhan Ye1, Zheyan Luo1, Zhangchi Feng1, Yongqiang Ma2 1School of Computer Science and Engineering, Beihang University, China 2School of Software and Microelectronics, Peking University, China {hiyouga,zhang.jh,yeyanhan,akamya,zcmuller}@buaa.edu.cn, zhangrc@act.buaa.edu.cn, codingma@pku.edu.cn Demonstration video: https://youtu.be/W29FgeZEpus
@@ 1 Introduction
Fine-tuning extremely large number of parameters with limited resources becomes the main challenge of adapting LLM to downstream tasks. A popular solution is efficient fine-tuning (Houlsby et al., 2019; Hu et al., 2022; Dettmers et al., 2023), which reduces the training cost of LLMs when adapting to various tasks. However, the community contributes various methods for efficient fine-tuning, lacking a systematic framework that adapts and unifies these methods to different LLMs and provides a friendly interface for user customization.
@@
LlamaFactory is open-sourced under the Apache-2.0 license. It has already garnered over 24,000 stars and 3,000 forks on the GitHub, and hundreds of open-source models have been built upon LlamaFactory on the Hugging Face Hub [^1] https://huggingface.co/models?other=llama-factory. For example, Truong et al. (2024) build GemSUra-7B based on LlamaFactory, revealing the cross-lingual abilities of Gemma (Mesnard et al., 2024). Furthermore, dozens of studies have utilized our framework to explore LLMs (Wang et al., 2023a; Yu et al., 2023; Bhardwaj et al., 2024).
2 Related Work
LlamaFactory FastChat LitGPT LMFlow Open-Instruct
LoRA
QLoRA
DoRA
LoRA+
PiSSA
GaLore
BAdam
Flash attention
S2 attention
Unsloth
DeepSpeed
SFT
RLHF
DPO
KTO
ORPO
Table 1: Comparison of features in LlamaFactory with popular frameworks of fine-tuning LLMs.
With the rapid increase in demand for fine-tuning LLMs, numerous frameworks for adapting LLMs to specific purposes have been developed. LLaMA-Adapter (Zhang et al., 2024) efficiently fine-tunes the Llama model (Touvron et al., 2023a) using a zero-initialized attention. FastChat (Zheng et al., 2023) is a framework focused on training and evaluating LLMs for chat completion purposes. LitGPT (AI, 2023) provides the implementation of generative models and supports various training methods. Open-Instruct (Wang et al., 2023d) provides recipes for training instruct models. Colossal AI (Li et al., 2023b) takes advanced parallelism strategies for distributed training. LMFlow (Diao et al., 2024) supports training LLMs for specialized domains or tasks. GPT4All (Anand et al., 2023) allows LLMs to run on consumer devices, while also providing fine-tuning capabilities. Compared with existing competitive frameworks, LlamaFactory supports a broader range of efficient fine-tuning techniques and training approaches. We list the features among representative frameworks in Table 1.
@@ 3 Efficient Fine-Tuning Techniques
Efficient LLM fine-tuning techniques can be divided into two main categories: those focused on optimization and those aimed at computation. The primary objective of efficient optimization techniques is to fine-tune the parameters of LLMs while keeping costs to a minimum. On the other hand, efficient computation methods seek to decrease the time or space for the required computation in LLMs. The methods included in LlamaFactory are listed in Table 2. We will present these efficient fine-tuning techniques and show the substantial efficiency improvement achieved by incorporating them into our framework in the following sections.
Freeze-tuning GaLore LoRA DoRA LoRA+ PiSSA
@@
Table 2: Compatibility between the fine-tuning techniques featured in LlamaFactory.
@@ 3.1 Efficient Optimization
Firstly, we provide an overview of the efficient optimization techniques utilized in LlamaFactory. The freeze-tuning method (Houlsby et al., 2019) involves freezing a majority of parameters while fine-tuning the remaining parameters in a small subset of decoder layers. Another method called gradient low-rank projection (GaLore) (Zhao et al., 2024) projects gradients into a lower-dimensional space, facilitating full-parameter learning in a memory-efficient manner. Similarly, BAdam (Luo et al., 2024) leverages block coordinate descent (BCD) to efficiently optimize the extensive parameters. On the contrary, the low-rank adaptation (LoRA) (Hu et al., 2022) method freezes all pre-trained weights and introduces a pair of trainable low-rank matrices to the designated layer. When combined with quantization, this approach is referred to as QLoRA (Dettmers et al., 2023), which additionally reduces the memory usage. DoRA (Liu et al., 2024) breaks down pre-trained weights into magnitude and direction components and updates directional components for enhanced performance. LoRA+ (Hayou et al., 2024) is proposed to overcome the sub-optimality of LoRA. PiSSA (Meng et al., 2024) initializes adapters with the principal components of the pre-trained weights for faster convergence.
@@ 3.2 Efficient Computation
In LlamaFactory, we integrate a range of techniques for efficient computation. Commonly utilized techniques encompass mixed precision training (Micikevicius et al., 2018) and activation checkpointing (Chen et al., 2016). Drawing insights from the examination of the input-output (IO) expenses of the attention layer, flash attention (Dao et al., 2022) introduces a hardware-friendly approach to enhance attention computation. S2 attention (Chen et al., 2024b) tackles the challenge of extended context with shifted sparse attention, thereby diminishing memory usage in fine-tuning long-context LLMs. Various quantization strategies (Dettmers et al., 2022a; Frantar et al., 2023; Lin et al., 2023; Egiazarian et al., 2024) decrease memory requirements in large language models (LLMs) by utilizing lower-precision representations for weights. Nevertheless, the fine-tuning of quantized models is restricted to the adapter-based techniques like LoRA (Hu et al., 2022). Unsloth (Han and Han, 2023) incorporates Triton (Tillet et al., 2019) for implementing the backward propagation of LoRA, which reduces floating-point operations (FLOPs) during gradient descent and leads to expedited LoRA training.
LlamaFactory seamlessly combines these techniques into a cohesive structure to enhance the efficiency of LLM fine-tuning. This results in a reduction of the memory footprint from 18 bytes per parameter during mixed precision training (Micikevicius et al., 2018) or 8 bytes per parameter in half precision training (Le Scao et al., 2022) to only 0.6 bytes per parameter. Further elaboration on the components in LlamaFactory will be provided in the subsequent section.
@@ 4 LlamaFactory Framework
LlamaFactory consists of three main modules: Model Loader, Data Worker, and Trainer. The Model Loader manipulates various model architectures for fine-tuning, supporting both large language models (LLMs) and vision language models (VLMs). The Data Worker processes data from different tasks through a well-designed pipeline, supporting both single-turn and multi-turn dialogues. The Trainer applies efficient fine-tuning techniques to different training approaches, supporting pre-training, instruction tuning and preference optimization. Beyond that, LlamaBoard provides a friendly visual interface to access these modules, enabling users to configure and launch individual LLM fine-tuning instance codelessly and monitor the training status synchronously. We illustrate the relationships between these modules and the overall architecture of LlamaFactory in Figure 1.
LlamaBoardExperiment ConfiguratorTraining Status MonitorTrainerOptimizationApproachesLoRAPiSSAGaLoreBAdamPre-trainSFTRLHFDPOModel LoaderData WorkerInitializationPatchesQuantizationAdaptersAligningLoadingPreprocessMergingPre-Trained ModelsConversational Datasets
@@ 4.1 Model Loader
This section initially presents the four components in Model Loader: model initialization, model patching, model quantization, and adapter attaching, followed by a description of our approach of adapting to a wide range of devices by handling the parameter float-point precision during fine-tuning.
@@
We utilize the Auto Classes of Transformers (Wolf et al., 2020) to load pre-trained models and initialize parameters. Specifically, we load the vision language models using the AutoModelForVision2Seq class while the rest are loaded using the AutoModelForCausalLM class. The tokenizer is loaded using the AutoTokenizer class along with the model. In cases where the vocabulary size of the tokenizer exceeds the capacity of the embedding layer, we resize the layer and initialize new parameters with noisy mean initialization. To determine the scaling factor for RoPE scaling (Chen et al., 2023), we compute it as the ratio of the maximum input sequence length to the context length of the model.
@@
To enable the S2 attention, we employ a monkey patch to replace the forward computation of models. However, we use the native class to enable flash attention as it has been widely supported since Transformers 4.34.0. To prevent excessive partitioning of the dynamic layers, we set the mixture-of-experts (MoE) blocks as leaf modules when we optimize the MoE models under DeepSpeed ZeRO stage-3 (Rasley et al., 2020).
@@
We automatically identify the appropriate layers to attach adapters through traversing the model layers. The low-rank adapters are attached to all the linear layers for a better convergence as suggested by (Dettmers et al., 2023). The PEFT (Mangrulkar et al., 2022) library provides an extremely convenient way to implement the adapter-based methods such as LoRA (Hu et al., 2022), rsLoRA (Kalajdzievski, 2023), DoRA (Liu et al., 2024) and PiSSA (Meng et al., 2024). We replace the backward computation with the one of Unsloth (Han and Han, 2023) to accelerate the training. To perform reinforcement learning from human feedback (RLHF), a value head layer is appended on the top of the transformer model, mapping the representation of each token to a scalar.
@@
We handle the floating-point precision of pre-trained models based on the capabilities of computing devices. For NVIDIA GPUs, we adopt bfloat16 precision if the computation capability is 8.0 or higher. Otherwise, float16 is adopted. Besides, we adopt float16 for Ascend NPUs and AMD GPUs and float32 for non-CUDA devices. In mixed precision training, we set all trainable parameters to float32 for training stability. Nevertheless, we retain the trainable parameters as bfloat16 in half precision training.
@@ 4.2 Data Worker
"system": "…", "tools": "…", "images": ["…"]}
Table 3: Dataset structures in LlamaFactory.
@@
LlamaFactory is designed for fine-tuning the text generative models, which is primarily used in chat completion. Chat template is a crucial component in these models, because it is highly related to the instruction-following abilities of these models. Therefore, we provide dozens of chat templates that can be automatically chosen according to the model type. We encode the sentence after applying the chat template using the tokenizer. By default, we only compute loss on the completions, while the prompts are disregarded (Taori et al., 2023). Optionally, we can utilize sequence packing (Krell et al., 2021) to reduce the training time, which is automatically enabled when performing generative pre-training.
@@ 4.3 Trainer
We integrate state-of-the-art efficient fine-tuning methods, including LoRA+ (Hayou et al., 2024), GaLore (Zhao et al., 2024) and BAdam (Luo et al., 2024) to the Trainer by replacing the default components. These fine-tuning methods are independent of the Trainer, making them easily applicable to various tasks. We utilize the trainers of Transformers (Wolf et al., 2020) for pre-training and SFT, while adopting the trainers of TRL (von Werra et al., 2020) for RLHF and DPO. We also include trainers of the advanced preference optimization methods such as KTO (Ethayarajh et al., 2024) and ORPO (Hong et al., 2024) from the TRL library. The tailored data collators are leveraged to differentiate trainers of various training approaches. To match the input format of the trainers for preference data, we build $2n$ samples in a batch where the first $n$ samples are chosen examples and the last $n$ samples are rejected examples.
@@
Allowing RLHF training on consumer devices is crucial for democratizing LLM fine-tuning. However, it is difficult because RLHF training requires four different models. To address this problem, we propose model-sharing RLHF, enabling entire RLHF training with no more than one pre-trained model. Concretely, we first train an adapter and a value head with the objective function for reward modeling, allowing the model to compute reward scores. Then we initialize another adapter and value head and train them with the PPO algorithm (Ouyang et al., 2022). The adapters and value heads are dynamically switched through the set_adapter and disable_adapter methods of PEFT (Mangrulkar et al., 2022) during training, allowing a single pre-trained model to serve as policy model, value model, reference model, and reward model simultaneously. To the best of our knowledge, this is the first method that supports RLHF training on consumer devices.
@@
We can combine the above trainers with DeepSpeed (Rasley et al., 2020; Ren et al., 2021) for distributed training. We adopt data parallelism to fully exploit the ability of computing devices. Leveraging the DeepSpeed ZeRO optimizer, the memory consumption can be further reduced via partitioning or offloading.
@@ 4.4 Utilities
Model Evaluation
We include several metrics for evaluating LLMs, including multiple-choice tasks such as MMLU (Hendrycks et al., 2021), CMMLU (Li et al., 2023a), and C-Eval (Huang et al., 2023), as well as calculating text similarity scores like BLEU-4 (Papineni et al., 2002) and ROUGE (Lin, 2004). This feature facilitates users to measure the abilities of the fine-tuned models.
@@ 4.5 LlamaBoard: A Unified Interface for LlamaFactory
Gemma-2B Llama2-7B Llama2-13B
Method Trainable Memory Throughput PPL Trainable Memory Throughput PPL Trainable Memory Throughput PPL
Params (GB) (Tokens/s) Params (GB) (Tokens/s) Params (GB) (Tokens/s)
Baseline / / / 11.83 / / / 7.53 / / / 6.66
Full-tuning 2.51B 17.06 3090.42 10.34 6.74B 38.72 1334.72 5.56 / / / /
Freeze-tuning 0.33B 8.10 5608.49 11.33 0.61B 15.69 2904.98 6.59 0.95B 29.02 1841.46 6.56
GaLore 2.51B 10.16 2483.05 10.38 6.74B 15.43 1583.77 5.88 13.02B 28.91 956.39 5.72
LoRA 0.16B 7.91 3521.05 10.19 0.32B 16.32 1954.07 5.81 0.50B 30.09 1468.19 5.75
QLoRA 0.16B 5.21 3158.59 10.46 0.32B 7.52 1579.16 5.91 0.50B 12.61 973.53 5.81
LlamaBoard is a unified user interface based on Gradio (Abid et al., 2019) that allows users to customize the fine-tuning of LLMs without writing any code. It offers a streamlined model fine-tuning and inference service, enabling users to easily explore the potential of LLMs in their environments. LlamaBoard has the following notable features.
Easy Configuration
LlamaBoard allows us to customize the fine-tuning arguments through interaction with the web interface. We provide default values for a majority of arguments that are recommended for most users, simplifying the configuration process. Moreover, users can preview the datasets on the web UI to check their correctness.
Monitorable Training
During the training process, the training logs and loss curves are visualized and updated in real time, allowing users to monitor the training progress. This feature provides valuable insights to analyze the fine-tuning process.
Flexible Evaluation
LlamaBoard supports calculating the text similarity scores on the datasets to automatically evaluate the model or performing human evaluation by chatting with the model.
Multilingual Support
LlamaBoard provides localization files, facilitating the integration of new languages for rendering the interface. Currently we support three languages: English, Russian and Chinese, which allows a broader range of users to utilize LlamaBoard for fine-tuning LLMs.